OpenStudy (anonymous):

find the derivative of y with respect to x, t, or theta; y=x^6lnx-1/3x^3

OpenStudy (oghorabah):

by putting " ln " for both sides we get \[\ln y = \ln x ^{(6\ln x - 1)\div 3x ^{3}}\] then we get this equation \[\ln y = ( ( 6\ln x - 1) \div 3x ^{3} ) \times \ln x\] this equation is long to derive but easier than its first form